OpenStudy (anonymous):

College Algebra: Reviewing for Final Exam! Expand the logarithmic expression: ln(3e^2) **Not sure how to solve this; not good with logarithms**

OpenStudy (anonymous):

@dan815 Think you can help me real quick?

OpenStudy (dan815):

big baby to the rescue

OpenStudy (anonymous):

:D Nice

OpenStudy (dan815):

ln(3e^2) okay ln(A*B)=lna+lnb

OpenStudy (dan815):

so ln(3*e^2)= ln3 +lne^2 =ln3+2

OpenStudy (anonymous):

Ohh, I get it! Oh god, thank you so much!! :) You're a big help! :)

OpenStudy (anonymous):

Quick question: Wouldn't the 2 come first in the equation?

OpenStudy (dan815):

sure u are welcome

OpenStudy (dan815):

it doesnt matter either way its equal

OpenStudy (dan815):

like if u add 2+3 or 3+2